Ethanol is a possible fuel.Use average bond energies to calculate ΔHrxn for the combustion of ethanol.

CH3CH2OH (g) + 3O2(g) ------> 2CO2 (g) +3H2O(g)

ΔHrxn= ______ kJ/mol
Answer #1
Concepts and reason

Write the balanced equation for the combustion of ethanol.

Calculate the value of enthalpy of reaction by using the standard values of bond energies of reactants and products.

Fundamentals

A balanced equation is an equation which contains all the reactants and products in a fixed ratio. A balanced equation has the following characteristics:

1)It should follow the law of conservation of mass.

2)The total number of different types of elements on reactant side should be equal to the product side.

3)Stoichiometric coefficients are put in front of the molecules on both the reactant and product side to balance a given reaction. The coefficients to be used are whole number or fractional number.

Combustion is a chemical process in which the substance is burned in the presence of oxygen. The products formed are called combustion products.

The amount of heat released on combustion of a substance is called the heat of combustion of that substance. It is denoted by ДН
. The unit of heat of combustion is kJ/mol
.

It is calculated by subtracting the bond energies of product from the reactants in a balanced reaction.

The combustion reaction of ethanol is written as follows:

CH, CH,OH +30, > 2C0, +3Н,0

The standard bond energies are as follows:

Bondenergy of C -C = 347 kJ /mol
Bondenergy of C-H 414KJ/mol
Bondenergy of C-0=360KJ/mol
Bondenergy of O-H 464kJ /mol

Bond energy of O = O = 498 kJ/mol
Bondenergy of C = O = 799 kJ /mol

The combustion reaction of ethanol is written as follows:

The enthalpy of the above reaction is the heat of combustion of ethanol.

The enthalpy of the above reaction is calculated as follows:

|AH
= Bond energy of reactants - Bond energy of products
=(5(cH)1(C-C)+1(C-0)+1(0-H)+3(0 0))-(6(0-H)+4(C=0))
-[(5(414)+1(347)

Ans:

The heat of combustion of ethanol, CH,CH,ОН
, per mole of ethanol is -1245 kJ/mol
.